OpenStudy (anonymous):

What is 0.514 written as a power of 10. is it 514x10^2?

OpenStudy (anonymous):

514*10^2 = 514*100=51400

OpenStudy (anonymous):

oh.. then isit 514 x 10^-3?

OpenStudy (anonymous):

thats one answer, scientific notation would be 5.14*10^-1

OpenStudy (anonymous):

one answer? so is it right?

OpenStudy (callisto):

0.1 = 1x10^-1 0.11 = 1.1 x 10^-1 Now, your question is 0.514, can you guess what the answer would be?

OpenStudy (anonymous):

i dont get it sorry

OpenStudy (callisto):

0.1 = 1x10^-1 <- get this one?

OpenStudy (anonymous):

ya

OpenStudy (callisto):

0.11 = 1.1 x 10^-1 <- How about this one?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

so do i put 514 in the 1.1 place?

OpenStudy (callisto):

Not really. Can you do this one first: 0.111 ?

OpenStudy (anonymous):

1.11 x 10^-1 ?

OpenStudy (callisto):

Yup! Now, you have 0.514 instead of 0.111. Can you do it??

OpenStudy (anonymous):

514 x 10^-1?

OpenStudy (callisto):

Nope ._.! For 0.111 , you got 1.11 x 10^-1 Now you have 0.514 , you'll get ____ x 10^-1 ?

OpenStudy (anonymous):

oh is it 5.14?

OpenStudy (callisto):

Yes!

OpenStudy (callisto):

for the blank...

OpenStudy (callisto):

Once again, what is 0.514?

OpenStudy (anonymous):

so is the answer 5.14 x 10^-1?

OpenStudy (callisto):

Yes!

OpenStudy (anonymous):

ok thank you so much. :)

OpenStudy (callisto):

You're welcome :)

OpenStudy (anonymous):

hey! are you doing odysseyware?